# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 2.2 Normalizing Flow in Function Space
[…]
The core of the model consists of two components: the pre-transformed measure and the transformations. The pre-transformed measure should be simple enough to enable efficient sampling, while the transformations must be flexible enough to convert this simple measure into a complex one that accurately approximates the target measure. We select the prior $\mu_{0}$ of Bayes’ formula (2.3) as the pre-transformed measure. The transformations are a series of parameterized operators, denoted as $f_{\theta_{n}}^{(n)}$, mapping from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$. Here, $n=1,2,\ldots,N$, and $\theta_{n}\in\Theta_{n}$ represents the parameters of $f_{\theta_{n}}^{(n)}$. The space $\Theta_{n}$ is a parameter space, whose specific form depends on the parameterization of $f_{\theta_{n}}^{(n)}$. For instance, in the context of functional planar flow illustrated in Subsection 2.4, the parameter space $\Theta_{n}$ is defined as follows:
@@
where $\mathcal{H}$ is the Cameron-Martin space of the prior $\mu_{0}$. Let $\theta=\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ denote the collection of parameters. The final complex measures are constructed by composing a series of transformations, which convert the pre-transformed measure into the post-transformed one. We write that
@@
and denote $u_{\theta}^{N}:=f_{\theta}(u_{0})$ where $u_{0}$ is a sample from $\mu_{0}$. Then $u_{\theta}^{N}$ will be a sample from the measure $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$, where $\mu_{0}\circ f_{\theta}^{-1}$ is the law of $f_{\theta}$ with respect to $\mu_{0}$ [7]. By varying the parameters $\theta$, the transformed measures $\mu_{f_{\theta}}$ form a set of measures on $\mathcal{H}_{u}$, denoted by
@@
Functional normalizing flow gives us a way to specify the approximate measures required for variational inference. However, since all the measures in $\mathcal{M}(\mathcal{H}_{u})$ should be equivalent to the prior $\mu_{0}$ [39], the choice of the parametric transformations $\{f_{\theta_{n}}^{(n)}\}_{n=1}^{N}$ may pose significant challenges. The measures within the infinite-dimensional function space are inherently prone to be singular with each others. To elucidate this concept, we present two pertinent examples:
@@
It is worth noting that although the transformations $f_{m}$ and $f_{2}$ are quite simple, they will convert $\mu_{0}$ into measures that are singular with respect to $\mu_{0}$, which contradicts the theoretical framework of infinite-dimensional variational inference. Consequently, the judicious selection of transformations is paramount for normalizing flow in function space, requiring both model flexibility and the preservation of measure equivalence. Here, we introduce a general theorem characterizing the transformations of functional normalizing flow. Assuming $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator mapping $\mathcal{H}_{u}$ to itself for $n=1,2,\ldots,N$, parameterized by $\theta_{n}$, $D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ represents the Fréchet derivative of $\mathcal{F}_{\theta_{n}}^{(n)}$, and $\theta=\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$. We define transformations as follows:
@@
We now present the following theorem, which establishes conditions ensuring the equivalence of measures before and after the transformations defined in equation (2.6).
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$. The composite transformation $f_{\theta}$ is given by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator.
  • 3.
Then we will know that $\mu_{f_{\theta}}$ is equivalent with $\mu_{0}$ (denoted by $\mu_{f_{\theta}}\sim\mu_{0}$).
Theorem 2.3 establishes the equivalence between transformed measures and the prior $\mu_{0}$. Consequently, the collection of transformed measures, denoted $\mathcal{M}(\mathcal{H}_{u})$, constitutes a suitable set of approximating measures for variational inference.
Analyzing the KL divergence between the approximate measure $\mu_{f_{\theta}}$ and the posterior $\mu$ necessitates the RN derivative of the transformed measure $\mu_{f_{\theta}}$ with respect to the prior $\mu_{0}$. We present a theorem that provides an expression for this derivative.
Theorem 2.4.
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$. The composite transformation $f_{\theta}$ is given by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
  • 1.
  • 2.
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator.
@@
Then the RN derivative of $\mu_{f_{\theta}}$ with respect to $\mu_{0}$ is given by:
@@
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for each $n=1,2,\ldots,N$, and ${\det}_{1}(Df_{\theta_{n}}^{n}(u_{n-1}))$ is the Fredholm-Carleman determinant [7] of the linear operator $Df_{\theta_{n}}^{n}(u_{n-1})$ (A rigorous definition can be found in the Appendix).
Leveraging Theorem 2.4, we can compute the necessary KL divergence
[…]
@@
Consequently, the optimization problem (2.5) can be reformulated as follows:
@@
Monte Carlo sampling allows us to estimate the KL divergence, which we can use as a loss function to optimize the transformation parameters $\theta$. The full details of the algorithm are presented in Algorithm 1. The estimated gradient of $\theta_{k}$ in Algorithm 1 can be computed by the preconditioned stochastic gradient-based optimization methods such as RMSprop (Root Mean Square Propagation), AdaGrad (Adaptive Gradient), and Adam (Adaptive Moment Estimation). In this article, we adopt the Adam optimizer, with implementation details provided in Section 4.
@@
1:  Initialize the hyperparameter $\lambda$ to $\lambda_{0}$, set the sampling size to $N$, the iterative number to $K$, and the learning rate schedule to $\alpha_{k}$. Then, initialize the iteration counter $k$ to 00;
@@
3:     Sampling $N$ functions $\{u_{i}\}_{i=1}^{N}$ from prior $\mu_{0}$;
@@ 2.3 Linear Transformation
Let $\mathcal{H}_{u}$ be a separable Hilbert space, $f_{\theta_{n}}^{(n)}:\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$, and $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is a bounded linear operator. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ is a finite rank operator.
@@ 2.3.1 Functional Householder Flow
The first model, termed functional Householder flow, is a generalization of the Householder flow in Euclidean space [35, 45]. Each layer of functional Householder flow employs the following transformation:
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u-0.5v_{n}(\langle v_{n},u\rangle_{\mathcal{H}_{u}}+b_{n})$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$ and $\theta_{n}=\{v_{n},b_{n}\}$ be its parameters. The composite transformation $f_{\theta}$ is given by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure, given by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. Assuming the following conditions hold for all $n=1,2,\ldots,N$ and any value of $\theta_{n}$:
@@
Then we have $\mu_{f_{\theta}}\sim\mu$ and
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for each $n=1,2,\ldots,N$.
The parameters $\{v_{n}\}_{n=1}^{N}$ of the flow model, being functions themselves, require an efficient parameterization. This is essential not only for computational efficiency but also for ensuring compliance with the constraints outlined in the theorem. A novel parameterization method for the flow model is proposed here. We denote $\{\phi_{i}\}_{i=1}^{M}$ as the first $M$ eigenvectors of the prior covariance operator. The parameterization process for $v_{n}$ involves two steps. Firstly, an auxiliary vector $\hat{v}_{n}$ is introduced and parameterized as

### v2

@@ 2.2 Normalizing Flow in Function Space
In this paper, we draw inspiration from the ideas presented in Euclidean space planarflow. A series of parametric transformations will be applied to a simple measure to generate a flexible set of complex measures. This collection of measures forms the approximating measures set $\mathcal{M}(\mathcal{H}_{u})$ for variational inference as outlined in (2.5). Consistent with normalizing flow in Euclidean space planarflow, sylvesterflow, householderflow1, we refer to this set of models as functional normalizing flow.
The core of the model consists of two components: the pre-transformed measure and the transformations. The pre-transformed measure should be analytically tractable to enable efficient sampling, while the transformations must be flexible to convert this simple measure into a complex one that accurately approximates the target measure. We select the prior $\mu_{0}$ of Bayes’ formula (2.3) as the pre-transformed measure. The transformations are a series of parameterized operators, denoted as $f_{\theta_{n}}^{(n)}$, mapping from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$. Here, $n=1,2,\ldots,N$, and $\theta_{n}\in\Theta_{n}$ represents the parameters of $f_{\theta_{n}}^{(n)}$. The space $\Theta_{n}$ is a parameter space, whose specific form depends on the parameterization of $f_{\theta_{n}}^{(n)}$. For instance, in the context of functional planar flow illustrated in Subsection 2.5, the parameter space $\Theta_{n}$ is defined as follows:
@@
where $\mathcal{H}$ is the Cameron-Martin space of the prior $\mu_{0}$. Let $\theta=\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ denote the collection of parameters. The post-transformed measures are constructed by composing a series of transformations, which convert the pre-transformed measure into a complex one. We write that
@@
and denote $u_{\theta}^{N}\mathrel{\mathop{\ordinarycolon}}=f_{\theta}(u_{0})$ where $u_{0}$ is a sample from $\mu_{0}$. Then $u_{\theta}^{N}$ will be a sample from the measure $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$, where $\mu_{0}\circ f_{\theta}^{-1}$ is the law of $f_{\theta}$ with respect to $\mu_{0}$ gaussianmeasure. By varying the parameters $\theta$, the transformed measures $\mu_{f_{\theta}}$ form a set of measures on $\mathcal{H}_{u}$, denoted by
@@
Functional normalizing flow gives us a way to specify the approximate measures required for variational inference. However, since all the measures in $\mathcal{M}(\mathcal{H}_{u})$ should be equivalent to the prior $\mu_{0}$ Pinski2015SIAMMA, the choice of the parametric transformations $\{f_{\theta_{n}}^{(n)}\}_{n=1}^{N}$ may pose significant challenges. The measures within the infinite-dimensional function space are inherently prone to be singular with each others. To elucidate this concept, we present two pertinent examples:
@@
It is noteworthy that although the transformations $f_{m}$ and $f_{2}$ are simple, they will convert $\mu_{0}$ into measures that are singular with respect to $\mu_{0}$, which contradicts the theoretical framework of infinite-dimensional variational inference. Consequently, the judicious selection of transformations is essential for normalizing flow in function space, requiring both model flexibility and the preservation of measure equivalence. Here, we introduce a general theorem characterizing the transformations of functional normalizing flow. Assuming $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator mapping $\mathcal{H}_{u}$ to itself for $n=1,2,\ldots,N$, parameterized by $\theta_{n}$, $D\mathcal{F}_{\theta_{n}}^{(n)}(u)$ represents the Fréchet derivative of $\mathcal{F}_{\theta_{n}}^{(n)}$, and $\theta=\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$. We define transformations as follows:
@@
We now present the following theorem, which establishes conditions ensuring the equivalence of measures before and after the transformations defined in equation (2.6), and provides their corresponding RN derivative.
@@
Let $\mathcal{H}_{u}$ be a separable Hilbert space equipped with the Gaussian measure $\mu_{0}=\mathcal{N}(0,\mathcal{C}_{0})$, and $\mathcal{H}=\mathcal{H}(\mu_{0})$ be the Cameron-Martin space of the measure $\mu_{0}$. For each $n=1,2,\ldots,N$, let $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is an operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{u}$. The composite transformation $f_{\theta}$ is given by $f_{\theta}(u)=f_{\theta_{N}}^{(N)}\circ f_{\theta_{N-1}}^{(N-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$, and $\mu_{f_{\theta}}$ denotes its corresponding push-forward measure by $\mu_{f_{\theta}}=\mu_{0}\circ f_{\theta}^{-1}$. For any $\theta\in\Theta$, assume that for all $n=1,2,\ldots,N$ the following conditions hold:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ has finite dimensional range.
@@
Then $\mu_{f_{\theta}}$ is equivelent with $\mu_{0}$, and the RN derivative of $\mu_{f_{\theta}}$ with respect to $\mu_{0}$ is given by:
@@
where $u_{n}=f_{\theta_{n}}^{(n)}\circ f_{\theta_{n-1}}^{(n-1)}\circ\cdots\circ f_{\theta_{1}}^{(1)}(u)$ for each $n=1,2,\ldots,N$, and ${\det}_{1}(Df_{\theta_{n}}^{(n)}(u_{n-1}))$ is the Fredholm-Carleman determinant gaussianmeasure of the linear operator $Df_{\theta_{n}}^{(n)}(u_{n-1})$ (A rigorous definition can be found in the Appendix).
Theorem 2.3 establishes the equivalence between transformed measures $\mu_{f_{\theta}}$ and the prior $\mu_{0}$. Consequently, the collection of transformed measures, denoted $\mathcal{M}(\mathcal{H}_{u})$, constitutes a suitable set of approximating measures for variational inference. Leveraging Theorem 2.3, we can compute the necessary KL divergence
$D_{\text{KL}}(\mu_{f_{\theta}}||\mu)$ $=\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu}(u)\right)d\mu_{f_{\theta}}(u)$
$=\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu_{0}}(u)\right)d\mu_{f_{\theta}}(u)-\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu}{d\mu_{0}}(u)\right)d\mu_{f_{\theta}}(u)$
@@
The optimization problem (2.5) thus yields the following reformulation:
@@
Monte Carlo sampling allows us to estimate the KL divergence, which serves as the loss function, to optimize the transformation parameters $\theta$. Details of the algorithm are presented in Algorithm 1. The estimated gradient of $\theta_{k}$ in Algorithm 1 can be computed by the preconditioned stochastic gradient-based optimization methods such as RMSprop (Root Mean Square Propagation) RMS, AdaGrad (Adaptive Gradient) Adagrad, and Adam (Adaptive Moment Estimation) Adam. In this article, we adopt the Adam optimizer, with implementation details provided in Section 4.
@@
1: Initialize the hyperparameter $\lambda$ to $\lambda_{0}$, parameter $\theta$ to $\theta_{0}$, set the sampling size to $N$, the iterative number to $K$, and the learning rate schedule to $\alpha_{k}$. Then, initialize the iteration counter $k$ to 0;
@@
3:  Sampling NN functions $\{u_{i}\}_{i=1}^{N}$ from prior $\mu_{0}$;
@@
Remark 2.1.
The loss function involves terms expressed in the $\mathcal{H}$-norm, such as $\langle f_{\theta}(u)-u,f_{\theta}(u)-u\rangle_{\mathcal{H}}$ and $\langle u,u-f_{\theta}(u)\rangle_{\mathcal{H}}$. Note that the $\mathcal{H}$-inner product is defined by
$\langle u,v\rangle_{\mathcal{H}}=\langle\mathcal{C}_{0}^{-\frac{1}{2}}u,\mathcal{C}_{0}^{-\frac{1}{2}}v\rangle_{{H}_{u}},$
where $\mathcal{C}_{0}$ is a trace class operator. Without appropriate regularity conditions on the arguments of the $\mathcal{H}$-norm terms, these terms are likely to be infinite. However, to satisfy the measure equivalence requirement of Theorem 2.3 between the approximate and prior measures, the arguments must possess sufficient regularity, which in turn necessitates the use of the $\mathcal{H}$-norm. In fact, the equivalence of the post-transformed measure $\mu_{f_{\theta}}$ with the prior $\mu_{0}$, combined with established results from stuart2010inverse, Cotter2009IP, typically ensures that the approximate measure is equivalent to the posterior. This equivalence is critical because it ensures the KL divergence (the loss function) is well-defined and finite.
2.3 Connection with conventional neural operator
Our proposed functional normalizing flow (FNF) shares a similar neural architecture with the conventional neural operator kovachki2023neural: Both of them are neural operators designed to map between function spaces. Upon training, they both accept input functions of arbitrary dimension and exhibits discrete invariance. However, FNF and neural operator address fundamentally different problems, leading to distinct theoretical limitations. We elaborate on their differences and connections in detail below to facilitate the reader’s understanding.
Different Target: The objective of the neural operator differs from that of the FNF. Specifically, the neural operator aims to learn an operator $\mathcal{G}$ that establishes a one-to-one correspondence between two function spaces. Given data pairs $\{u_{i},\mathcal{G}(u_{i})\}_{i=1}^{N}$, the goal is to approximate $\mathcal{G}$ by a neural network $f_{\theta}$; a canonical example is when $\mathcal{G}$ is the solution operator of a PDE, mapping a source term $u_{i}$ to its corresponding solution $\mathcal{G}(u_{i})$.
In contrast, FNF aims to construct a mapping $f_{\theta}$ that transforms a tractable source measure $\mu_{0}$ (prior) into a complex target measure $\mu$ (posterior). Specifically, given a set of samples $\{u_{i}\}_{i=1}^{N}$ drawn from $\mu_{0}$, their images under $f_{\theta}$ form the set $\{f_{\theta}(u_{i})\}_{i=1}^{N}$. This resulting set consists of samples from the flow-induced measure $\mu_{f_{\theta}}$, which approximates samples from $\mu$.
Through training, the neural operator approach focuses on whether the prediction $f_{\theta}(u_{i})$ closely matches the ground-truth $\mathcal{G}(u_{i})$ for each input function $u_{i}$. In contrast, FNF shifts focus away from the individual relationship between $u_{i}$ and $f_{\theta}(u_{i})$, directing attention instead to the entire function ensembles $\{u_{i}\}_{i=1}^{N}$ and $\{f_{\theta}(u_{i})\}_{i=1}^{N}$. Its primary objective is to verify whether $\{f_{\theta}(u_{i})\}_{i=1}^{N}$ can be regarded as approximate samples from the target measure $\mu$ when the input ensemble $\{u_{i}\}_{i=1}^{N}\sim\mu_{0}$.
Different Loss Function: During training, the neural operator uses a dataset of paired samples $\{u_{i},\mathcal{G}(u_{i})\}_{i=1}^{N}$. Its loss function is typically defined as
$L(\theta)=\sum_{i=1}^{N}d\big(\mathcal{G}(u_{i}),f_{\theta}(u_{i})\big),$
where $d$ denotes a metric quantifying the distance between functions.
The loss function for FNF is designed to measure the distance between two probability measures $\mu_{f_{\theta}}$ and $\mu$. In our work, we adopt the KL divergence $D_{\text{KL}}(\mu_{f_{\theta}}||\mu)$ as the loss function, which can be write as
$D_{\text{KL}}(\mu_{f_{\theta}}||\mu)$ $=\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu}(u)\right)d\mu_{f_{\theta}}(u)$
$=\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu_{0}}(u)\right)d\mu_{f_{\theta}}(u)-\int_{\mathcal{H}_{u}}\ln\left(\frac{d\mu}{d\mu_{0}}(u)\right)d\mu_{f_{\theta}}(u)$
$=E_{\mu_{f_{\theta}}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu_{0}}(u)\right)-E_{\mu_{f_{\theta}}}\ln\left(\frac{d\mu}{d\mu_{0}}(u)\right)$
$=E_{\mu_{0}}\ln\left(\frac{d\mu_{f_{\theta}}}{d\mu_{0}}(f_{\theta}(u))\right)-E_{\mu_{0}}\ln\left(\frac{d\mu}{d\mu_{0}}(f_{\theta}(u))\right).$
It can be seen that the training procedure of the proposed algorithm only requires samples $\{u_{i}\}_{i=1}^{N}\sim\mu_{0}$ and does not rely on a large dataset of paired input-output instances $\{u_{i},\mathcal{G}(u_{i})\}_{i=1}^{N}$, contrasting with the dataset required by neural operator.
Distinct Theoretical Requirements: In contrast to finite-dimensional settings, probability measures defined on infinite-dimensional function spaces are inherently more prone to mutual singularity. As shown in previous studies on infinite-dimensional variational inference pinski2015algorithms, Pinski2015SIAMMA, jia2021stein, Jia2022VINet, jia2021variational, preserving the equivalence between the approximate measure $\mu_{f_{\theta}}$ and the pre-transformed measure $\mu_{0}$ is essential. This requirement imposes additional theoretical constraints on our model relative to neural operator, with precise mathematical formulations provided in Theorem 2.3.
@@ 2.4 Linear Transformation
Let $\mathcal{H}_{u}$ be a separable Hilbert space, $f_{\theta_{n}}^{(n)}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathcal{H}_{u}$, and $f_{\theta_{n}}^{(n)}(u)=u+\mathcal{F}_{\theta_{n}}^{(n)}(u)$, where $\mathcal{F}_{\theta_{n}}^{(n)}$ is a bounded linear operator. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
    The operator $\mathcal{F}_{\theta_{n}}^{(n)}$ has finite dimensional range.
@@ 2.4.1 Functional Householder Flow
The first model, termed functional Householder flow, is a generalization of the Householder flow in Euclidean space householderflow1, Householderflow2. Each layer of functional Householder flow employs the following transformation:
@@
Let $f_{\theta_{n}}^{(n)}(u)=u-0.5v_{n}(\langle v_{n},u\rangle_{\mathcal{H}_{u}}+b_{n})$. Assume that the following conditions hold for all $n=1,2,\ldots,N$:
@@
Then $f_{\theta_{n}}^{(n)}$ satisfies all the conditions of Theorem 2.3
The parameters $\{v_{n}\}_{n=1}^{N}$ of the flow model, being functions themselves, require an efficient parameterization. This is essential not only for computational efficiency but also for ensuring compliance with the constraints outlined in the theorem. A novel parameterization method for the flow model is proposed here. We denote $\{\phi_{i}\}_{i=1}^{M}$ as the first $M$ eigenfunctions of the prior covariance operator. The parameterization process for $v_{n}$ involves two steps. Firstly, an auxiliary function $\hat{v}_{n}$ is introduced and parameterized as
@@
Remark 2.2.
Similar with article chen2024learning, we first require the eigenfunctions of the covariance operator $\mathcal{C}_{0}$ before running the algorithm. While these eigenfunctions can sometimes be derived analytically, numerical methods are necessary when analytical solutions are infeasible. For instance, we can discretize the operator $\mathcal{C}_{0}$ using the method in bui2013computational and numerically obtain the eigenfunctions at some discrete level, employing techniques such as the double-pass algorithm saibaba2016randomized or any other classical algorithm presented in golub2013matrix.